# Improving Data Curation of Software Vulnerability Patches through Uncertainty Quantification

### v1

@@ III-E2 Separating Epistemic from Aleatoric Uncertainty
2Function DataCurationWithEHAL ($\mathbb{D}_{seed}$, $\mathbb{X}_{pool}$, $n$)
3       $\mathbf{D}_{new}$leftarrow DataSelection ($\mathbb{D}_{seed}$, $\mathbb{X}_{pool}$, EHALn)
4       return $\mathbf{D}_{new}$
5
6
Input: $\mathbb{X}_{pool}$: pool of patches to select from
   $\mathbb{U}_{pool}$: patches in $\mathbb{X}_{pool}$ with uncertainties
   $n$: $n$ patches to select
Output: $\mathbf{D}$: list of selected patches
7 Function EHALn ($\mathbb{X}_{pool}$, $\mathbb{U}_{pool}$, $n$)
8       $\mathbf{D}\leftarrow[\leavevmode\nobreak\ ]$
9       while $\mathbb{U}_{pool}\neq\emptyset\land\mathbf{D}.size()<n$ do
10             $d$leftarrow EHAL ($\mathbb{X}_{pool}$, $\mathbb{U}_{pool}$)
11             $\mathbb{X}_{pool}\leftarrow\mathbb{X}_{pool}\setminus\{d\}$
12             $\mathbb{U}_{pool}\leftarrow\mathbb{U}_{pool}\setminus\{d\}$
13             $\mathbf{D}.append(d)$
14
15      return $\mathbf{D}$
16
17
Input: $\mathbb{X}_{pool}$: pool of patches to select from
   $\mathbb{U}_{pool}$: patches in $\mathbb{X}_{pool}$ with uncertainties
Output: $d_{epi}$: selected patch
18
19Function EHAL ($\mathbb{X}_{pool}$, $\mathbb{U}_{pool}$)
20       $\mathbb{X}_{candidate}\leftarrow\mathbb{X}_{pool}$
21       $d_{epi}\leftarrow\max_{epi}(\mathbb{X}_{candidate},\mathbb{U}_{pool})$
22       $\mathbb{D}_{ali}\leftarrow\max_{ale}(\mathbb{X}_{candidate},\mathbb{U}_{pool},n_{ali})$
23       while $d_{epi}\in\mathbb{D}_{ali}$ do
24             $\mathbb{X}_{candidate}\leftarrow\mathbb{X}_{candidate}\setminus\{d_{epi}\}$
25             $d_{epi}\leftarrow\max_{epi}(\mathbb{X}_{candidate})$
26             $\mathbb{D}_{ali}\leftarrow\max_{ale}(\mathbb{X}_{candidate},n_{ali})$
27
28      return $d_{epi}$
@@ 26             $\mathbb{D}_{ali}\leftarrow\max_{ale}(\mathbb{X}_{candidate},n_{ali})$
Given an accurate estimation of epistemic and aleatoric uncertainty for patch data curation, as described in RQ1, it is still an open problem of how we use these two quantities to improve patch data curation. In this RQ2, we focus on the practical aspects of using UQ in ML-based patch data curation in order to obtain higher quality datasets. To this end, we design the EHAL Heuristic (Algorithm 1) that actively and selectively adds patches based primarily on high epistemic uncertainty to include in the dataset.
Instead of leveraging a formula that combines both epistemic and aleatoric uncertainties, the EHAL heuristic takes a “select-and-then-reject” approach. As shown in Algorithm 1, it first identifies patch instances with the highest epistemic uncertainty as candidates at Line 1, but rejects them if they are among the patches with the highest aleatoric uncertainty at Lines 1 to 1, i.e., it rejects those on the lowest extreme of the quality spectrum. The name of the heuristic is drawn from this idea as well: Epistemic High Aleatoric Low (EHAL).
To evaluate the EHAL Heuristic, we consider two baselines. One is the opposite to the EHAL Heuristic, i.e., Epistemic Low Aleatoric High (ELAH). ELAH selects the patches with the lowest epistemic uncertainty, and rejects them if they are also among those with the lowest aleatoric uncertainty. In addition, we also compare our algorithm with the random baseline that selects patches randomly.
In our evaluation, we gradually expand the training dataset, using one of the three algorithms: the proposed one and the two baseline algorithms. For each, we divide the training data partition randomly into two parts: the initial training dataset (20% of available patches) and the candidate training instances pool (60% of available patches). We construct the training dataset, initially the initial training dataset, and train a patch identification model. With the trained model, we estimate uncertainty measures, such as aleatoric and epistemic uncertainties. Based on the uncertainties, we select 10% of instances from the candidate pool and add these to the training dataset. We then retrain the model and repeat the process. We stop when we exhaust the candidate dataset. To evaluate the model’s predictive performance, we compute the F1 score on the test dataset (20% of available patches) as the size of the training dataset increases. To ensure the reliability of the results, we repeat the process 10 times, each beginning with a random shuffling of the entire dataset.
@@
The results for both datasets show that the model using EHAL performs the best (F1 score) as patches are selected and added to the training dataset. The model with the ELAH baseline performs the worst, while random selection is in the middle. This serves as evidence that the proposed algorithm can effectively select patches and improve predictive performance. Note, however, that as the percentage of added training data approaches 100%, there is less leeway for selection, i.e., the dataset runs out of ”good” candidates, and adding more patches to the training dataset has no discernible effect.
@@
IV-A Evaluating Patch Curation’s Impact on Automated Software Vulnerability Prediction
@@
    Train the patch curation model. We train a patch curation model using the VCMatch dataset and select the model that demonstrates the highest patch identification performance within the VCMatch dataset, leveraging 40% of the training instances (see Figure 5).
@@
    Extract commit patches from the Big-Vul dataset. We identify commit hashes associated with vulnerable functions in Big-Vul and use these hashes to extract the relevant patches.
@@
  5. 5.
    Train and evaluate LineVul. In the Big-Vul data, a vulnerable function has a number of correspondent non-vulnerable functions, all of these are indexed by the commit id of the security patch from which the vulnerable function is extracted. This organization of the Big-Vul data allows us to assemble the training data from the selected security patches. We train a LineVul model from scratch using the functions indexed by the commit id of the selected patches, and evaluate the performance of the trained LineVul model on the test partition. The vulnerable functions in the test partition is about 6.2%, and following the LineVul study [57], we report F1 score as the performance metric.
First, we examine the relationship between LineVul’s predictive performance and the amount of training data used. As a baseline, we run the identical experiments using a random data selection method. Figure 6 summarizes the experimental results where we show F1 score versus the amount of training data used. LineVul’s predictive performance varies given different test data. The figure exhibits the variance using a standard boxplots showing min, first quartile, median, third quartile, and max. The figure shows that (a) when we use the proposed patch selection method, LineVul achieves a higher F1 score with 40-80% of the data than with 100% of the training data, and (b) the proposed method consistently outperforms random selection except the case when nearly 90% or more patches are selected.
@@
Additionally, in Table V we compare the mean F1 scores obtained when using 100% of training data, using the training data selected by the proposed data curation method, and using random selection. When comparing the proposed data curation method with no data curation, i.e, using 100% of training data, the improvement of mean F1 score ranges from 1.47% to 2.02%. We observe a sightly higher improvement of F1 score when comparing the proposed data selection with the random data selection.
TABLE V: Comparing Mean F1 Scores: using 100% Training Data (F1 Score=0.834), using Proposed Data Curation Method, and Random Selection
F1 Score Gain
SelectedCuration Random vs. Random vs. 100%
40% 0.827 0.847 2.51% 1.61%
50% 0.826 0.847 2.55% 1.57%
60% 0.832 0.851 2.21% 2.02%
70% 0.835 0.849 1.66% 1.76%
80% 0.832 0.846 1.67% 1.47%
@@
SelectedCuration vs. 100% Curation vs. Random
40% $052.17e-$ 05 $065.41e-$ 06
50% $052.17e-$ 05 $065.41e-$ 06
60% $051.08e-$ 05 $047.52e-$ 04
70% $053.79e-$ 05 $042.44e-$ 04
80% $045.25e-$ 04 $041.62e-$ 04
@@
SelectedF1 GainTime Reduction
100% 0.834 —- 13h26m —-
50% 0.847 $1.61\%$ 5h53m $\frac{\text{13h26m}-\text{5h53m}}{\text{13h26m}}\approx 56\%$
The experimental result in Table VII shows that the proposed patch curation algorithm, which selects 40% of the training data, can benefit LineVul by gaining a small improvement in predictive performance of 1.61%. However, at the same time, the training time is substantially reduced by approximately 56%, i.e., a significant energy saving in training.
@@ V-A Vulnerability Patch Identification
In this work, we make two essential contributions. First, we provide a principled approach to assess the quality of data curation [36, 37, 38, 3, 39, 4]. Second, we offer guidance on data curation, focusing on not only data quality but also data usefulness, through the application of Uncertainty Quantification (UQ) techniques [13, 14]. As such, our work is not simply to “clean” existing datasets, rather it is aimed to curate highly useful security patches for improving down-stream applications of the patches.
@@ VI Threats to Validity
Generalization of Findings. Our research findings strongly suggest that Uncertainty Quantification (UQ) can serve as a systematic and principled approach to curate software changeset data, enabling effective vulnerability patch identification and other software quality assurance tasks. However, it is important to acknowledge a potential internal threat to the validity of these conclusions, as the generalizability of our results might be limited due to the number of models, datasets, and UQ approaches explored in our initial experiments. To address this internal validity concern and determine whether the observed results are merely coincidental, we conducted extensive array of experiments. By exploring various combinations of UQ techniques, datasets, and data features, we sought to ensure robustness in our findings. The key conclusions consistently emerged across all of our experiments, reinforcing the reliability of our main research outcomes.
@@ VII Conclusion and Future Work
Machine learning techniques that leverage historical vulnerabilities have become an important direction in software security assurance. However, obtaining high-quality vulnerability patch data, which is crucial for advancing this research, poses significant challenges. Several studies confirm substantial quality issues in software vulnerability patches collected from the NVD [13, 15]. Furthermore, there is a lack of a systematic approach to curate patch datasets and assess patch quality. Relying solely on software security experts for manual curation is impractical due to cost and limited availability of expertise. In addition, prior works primarily focus on data quality alone [15]. As machine learning techniques become an important research direction software security assurance, whether curated patches are informative to machine learning should be an important dimension to examine. However, patch data curation eyeing both utility values and quality is particular difficult to design due to the fact that both quality and usefulness span a spectrum.
@@
To address these concerns and challenges, our work aims to develop an approach for curating software changeset data for security software assurance. We propose an automatic, machine learning-based security patch curation approach intended to select patches with high quality and utility value. This is made possible by developing a heuristic that leverages both epistemic and aleatoric uncertainties computed on security patches. From our evaluation, we observe the following: 1) Model Ensemble is an effective method for producing reliable Uncertainty Quantification (UQ) in security patch curation; 2) software changeset data can exhibit different distributions across projects or components, making heteroscedastic modeling a superior approach that demonstrates better predictive performance and UQ quality compared to homoscedastic modeling; 3) epistermic uncertainty can serve as a proxy for the informativeness of the patches while aleatoric uncertainty can serve as a guidance for the quality of the patches; 4) our study highlights that UQ measures can guide the selection of changesets to benefit downstream secure software quality assurance tasks, such as a software vulnerability prediction model. For a manually curated high-quality dataset, the benefits come 1) a small but statistically significant improvement of predictive performance, and 2) a significant reduction in software vulnerability model training time (i.e., significant energy saving). Our work is a first to curate security patches with a focus on their utility values in addition to their quality, which sets the stage for the future work that aims to scale our experiments to other, larger vulnerability datasets and curate vulnerability patches from databases such as the NVD to significantly improve predictive performance and reduce uncertainty for machine learning-based software security assurance approaches.

### v2

@@ IV-A Automatic Curation of Vulnerability Patches with UQ
2Function DataCurationWithEHAL($\mathbb{X}_{\text{pool}}$, $\mathbb{U}_{\text{pool}}$, $n$, $n_{\text{ale}}$)
3  $\mathbb{D}_{\text{new}}\leftarrow[\penalty 10000\ ]$
4  while $\mathbb{U}_{\text{pool}}\neq\emptyset$ and $\mathbb{D}_{\text{new}}.\text{size}()<n$ do
5     $d_{\text{epi}}\leftarrow\textnormal{{EHAL}}(\mathbb{X}_{\text{pool}},\mathbb{U}_{\text{pool}},n_{\text{ale}})$
6     $\mathbb{X}_{\text{pool}}\leftarrow\mathbb{X}_{\text{pool}}\setminus\{d_{\text{epi}}\}$
7     $\mathbb{U}_{\text{pool}}\leftarrow\mathbb{U}_{\text{pool}}\setminus\{d_{\text{epi}}\}$
8     $\mathbb{D}_{\text{new}}.\text{append}(d_{\text{epi}})$
9
10 return $\mathbb{D}_{\text{new}}$
11
12Function EHAL($\mathbb{X}_{\text{pool}}$, $\mathbb{U}_{\text{pool}}$, $n_{\text{ale}}$ )
13  $\mathbb{X}_{\text{candidate}}\leftarrow\mathbb{X}_{\text{pool}}$
14  $\mathbb{U}_{\text{candidate}}\leftarrow\mathbb{U}_{\text{pool}}$
15  $d_{\text{epi}}\leftarrow\text{TopOneByEpistemicUQ}(\mathbb{X}_{\text{candidate}},\mathbb{U}_{\text{candidate}})$
16  $\mathbb{D}_{\text{ale}}\leftarrow\text{TopNByAleatoricUQ}(\mathbb{X}_{\text{candidate}},\mathbb{U}_{\text{candidate}},n_{\text{ale}})$
17  while $d_{\text{epi}}\in\mathbb{D}_{\text{ale}}$ do
18     $\mathbb{X}_{\text{candidate}}\leftarrow\mathbb{X}_{\text{candidate}}\setminus\{d_{\text{epi}}\}$
19     $\mathbb{U}_{\text{candidate}}\leftarrow\mathbb{U}_{\text{candidate}}\setminus\{d_{\text{epi}}\}$
20     $d_{\text{epi}}\leftarrow\text{TopOneByEpistemicUQ}(\mathbb{X}_{\text{candidate}},\mathbb{U}_{\text{candidate}})$
21     $\mathbb{D}_{\text{ale}}\leftarrow\text{TopNByAleatoricUQ}(\mathbb{X}_{\text{candidate}},\mathbb{U}_{\text{candidate}},n_{\text{ale}})$
22
23 return $d_{\text{epi}}$
@@ 19     $\mathbb{U}_{\text{candidate}}\leftarrow\mathbb{U}_{\text{candidate}}\setminus\{d_{\text{epi}}\}$
The results for both datasets show that the model using EHAL performs the best (F1 score) as patches are selected and added to the training dataset. The model with the ELAH baseline performs the worst, while random selection is in the middle. This serves as evidence that the proposed algorithm can effectively select patches and improve predictive performance. Note, however, that as the percentage of added training data approaches 100%, there is less leeway for selection, i.e., the dataset runs out of ”good” candidates, and adding more patches to the training dataset has no discernible effect. The modest gains observed in Figure 6 are partially due to the limitations of the SAP and VCMatch datasets, which contain only 1,200 to 1,700 instances. With datasets of this size, it is challenging to observe a large effect.
@@
The training data ratio is an important hyperparameter in this approach, as it influences the quality of the data curation model. Its optimal value varies depending on the dataset and the model. Our experiments show that, for the VCMatch dataset, performance gains plateau after using approximately 40% of the curated data, whereas for the higher-quality SAP dataset, 80% was needed to reach maximum utility. This variability underscores the importance of tailoring the method to specific use cases, since the balance between data quality, model performance, and training-time efficiency differs across real-world applications. The data curation process is inherently iterative. A practical approach is to start with a small random sample of labeled patches, estimate a suitable ratio, and then adjust it based on model performance as more data are added.
@@
IV-B Evaluating Patch Curation’s Impact on Automated Software Vulnerability Prediction
@@
    Train the patch curation model. We train a patch curation model using the VCMatch dataset and select the model that demonstrates the highest patch identification performance within the VCMatch dataset, leveraging 40% of the training instances (see Figure 6). This is because both the VCMatch dataset and the Big-Vul dataset consist of C/C++ code, while the SAP dataset contains solely Java code.
@@
    Remove overlapping projects from the Big-Vul dataset. The Big-Vul dataset contains functions from 310 software projects, among which 8 are also in the VCMatch dataset. To eliminate the risk of data contamination, we remove these 8 software projects, namely, FFmpeg, QEMU, OpenSSL, WireShark, PHP-SRC, Moodle, ImageMagick, and Linux, from the Big-Vul dataset completely. As a result, the number of functions in the Big-Vul dataset is reduced from 188,636 to 130,449. For convenience, in the remaining discussion of this section, we refer to this filtered version as the Big-Vul dataset.
@@
    Extract commit patches from the filtered Big-Vul dataset. We identify commit hashes associated with vulnerable functions in Big-Vul and use these hashes to extract the relevant patches.
  4. 4.
@@
  6. 6.
    Train and evaluate LineVul. In the Big-Vul data, a vulnerable function has a number of correspondent non-vulnerable functions, all of these are indexed by the commit id of the security patch from which the vulnerable function is extracted. This organization of the Big-Vul data allows us to assemble the training data from the selected security patches. We train a LineVul model from scratch using the functions indexed by the commit id of the selected patches, and evaluate the performance of the trained LineVul model on the test partition. The vulnerable functions in the test partition is about 6.1%, and following the LineVul study [65], we report F1 score as the performance metric.
First, we examine the relationship between LineVul’s predictive performance and the amount of training data used. As a baseline, we run the identical experiments using a random data selection method. Figure 7 summarizes the experimental results where we show F1 score versus the amount of training data used. LineVul’s predictive performance varies given different test data. The figure exhibits the variance using a standard boxplots showing min, first quartile, median, third quartile, and max. The figure shows that (a) when we use the proposed patch selection method, LineVul achieves a higher F1 score with 20-60% of the data than with 100% of the training data, and (b) the proposed method consistently outperforms random selection except the case when nearly 90% or more patches are selected.
@@
Additionally, in Table V we compare the mean F1 scores obtained when using 100% of training data, using the training data selected by the proposed data curation method, and using random selection. When comparing the proposed data curation method with no data curation, i.e, using 100% of training data, the improvement of mean F1 score ranges from 1.66% to 3.52%.We observe a higher improvement of F1 score when comparing the proposed data selection with the random data selection.
TABLE V: Comparing Mean F1 Scores: using 100% Training Data (Mean F1 Score=0.843), using Proposed Data Curation Method, and Random Selection
Mean F1 Score Gain
Selected % Random Curation vs. Random vs. 100%
20% 0.800 0.857 7.07% 1.66%
30% 0.818 0.868 6.02% 2.94%
40% 0.823 0.872 6.05% 3.52%
50% 0.825 0.858 3.98% 1.81%
60% 0.825 0.860 4.22% 2.08%
@@
Selected % Curation vs. Random Curation vs. 100%
20% 8.23e058.23e-05 1.43e241.43e-24
30% 2.68e042.68e-04 1.63e061.63e-06
40% 5.95e065.95e-06 1.84e071.84e-07
50% 6.15e046.15e-04 2.27e032.27e-03
60% 4.45e034.45e-03 5.17e035.17e-03
@@
Selected % F1 Gain % Time Reduction
100% 0.843 —- 8h57m —-
40% 0.872 $3.52\%$ 3h55m $\frac{\text{8h57m}-\text{3h55m}}{\text{8h57m}}\approx 56\%$
The experimental result in Table VII shows that the proposed patch curation algorithm, which selects 40% of the training data, can benefit LineVul by gaining a small improvement in predictive performance of 3.52% (Table VI). However, at the same time, the training time is substantially reduced by approximately 56%, i.e., a significant energy saving in training. This benefit is particularly valuable for large vulnerability data collections (that may even include data augmentation), which are increasingly important to modern ML models.
@@ V-A Vulnerability Patch Identification
In this work, we make two essential contributions. First, we provide a principled approach to assess the quality of data curation [34, 35, 36, 3, 37, 4]. Second, we offer guidance on data curation, focusing on not only data quality but also data usefulness, through the application of Uncertainty Quantification (UQ) techniques [13, 14]. As such, our work is not simply to “clean” existing datasets, rather it is aimed to curate highly useful security patches for improving down-stream applications of the patches. While dataset cleaning primarily addresses technical errors and inconsistencies, curation involves selecting data points that maximize utility, ensuring the resulting dataset is optimized for improving model performance and training efficiency.
@@ VI Threats to Validity
Generalization of Findings. Our research findings strongly suggest that Uncertainty Quantification (UQ) can serve as a systematic and principled approach to curate software changeset data, enabling effective vulnerability patch identification and other software quality assurance tasks. However, it is important to acknowledge a potential internal threat to the validity of these conclusions, as the generalizability of our results might be limited due to the number of models, datasets, and UQ approaches explored in our initial experiments. Furthermore, we acknowledge that our technique has thus far been applied only to the SAP, VCMatch, and Big-Vul datasets. To mitigate this internal validity concern, we conducted an extensive array of experiments. By exploring various combinations of UQ techniques, datasets, and data features, we sought to ensure robustness in our findings. The key conclusions consistently emerged across all of our experiments, reinforcing the reliability of our main research outcomes.
This study randomly splits labeled commit patches into training, validation, and test partitions, a standard practice in software vulnerability prediction research. While data leakage is a general concern in machine learning, random splits cannot fully prevent it due to potential dependencies between commits, such as change couplings and genealogies [71, 72]. In our case, this risk is minimal. The datasets consist of vulnerability patches and randomly selected clean patches, without full evolutionary histories. As a result, the chance of leakage from commit dependencies is low.
Sources of Uncertainty. A recent study calls into the attention of the effectiveness of noisy label learning on deep learning for program understanding [73]. The study indicates that deep learning models still struggle to detect real-world noise in program understanding datasets. In this research, we do not differentiate the sources of uncertainty, i.e., whether they stem from the feature space or the labeling space. A future direction will be to investigate whether UQ can improve noise label learning for program understanding by more accurately detecting labeling errors.
Programming Languages and Data Selection. Programming Languages and Data Selection. In this paper, we evaluated the impact of the proposed data selection approach on software vulnerability prediction using C/C++ code. Whether this data selection approach has a similar impact on software vulnerability prediction for other programming languages or in a cross-language setting remains an open question for future exploration.
@@ VII Conclusion and Future Work
Machine learning techniques that leverage historical vulnerabilities have become an important direction in software security assurance. However, obtaining high-quality vulnerability patch data, which is crucial for advancing this research, poses significant challenges. Several studies confirm substantial quality issues in software vulnerability patches collected from the NVD [13, 15]. Furthermore, there is a lack of a systematic approach to curate patch datasets and assess patch quality. Relying solely on software security experts for manual curation is impractical due to cost and limited availability of expertise. In addition, prior works primarily focus on data quality alone [15]. As machine learning techniques become an important research direction software security assurance, whether curated patches are informative to machine learning should be an important dimension to examine. However, designing patch data curation approaches that account for both utility and quality is particularly challenging to design due to the fact that both quality and usefulness span a spectrum.
@@
To address these concerns and challenges, our work aims to develop an approach for curating software changeset data for security software assurance. We propose an automatic, machine learning-based security patch curation approach intended to select patches with high quality and utility value. This is made possible by developing a heuristic that leverages both epistemic and aleatoric uncertainties computed on security patches. From our evaluation, we observe the following: 1) Model Ensemble is an effective method for producing reliable Uncertainty Quantification (UQ) in security patch curation; 2) software changeset data can exhibit different distributions across projects or components, making heteroscedastic modeling a superior approach that demonstrates better predictive performance and UQ quality compared to homoscedastic modeling; 3) epistemic uncertainty can serve as a proxy for the informativeness of the patches while aleatoric uncertainty can serve as a guidance for the quality of the patches; 4) in all evaluations, we compare data curated using the proposed approach with 100% of the data from VCMatch and Big-Vul, representing heuristic-based and machine learning-based state-of-the-art curation methods, respectively, and find that it selects higher-quality and more informative patches. Our study highlights that UQ measures can guide the selection of changesets to benefit downstream secure software quality assurance tasks, such as a software vulnerability prediction model. For a manually curated high-quality dataset, the benefits come 1) a small but statistically significant improvement of predictive performance, and 2) a significant reduction in software vulnerability model training time (i.e., significant energy saving). Our work is a first to curate security patches with a focus on their utility values in addition to their quality, which sets the stage for the future work that aims to scale our experiments to other, larger vulnerability datasets and curate vulnerability patches from databases such as the NVD to significantly improve predictive performance and reduce uncertainty for machine learning-based software security assurance approaches.
Future work also includes assessing new UQ models aimed at enabling a data curation pipeline that further improves downstream applications. Additionally, advances in large language models (LLMs) have led to the exploration of non-training methods, such as in-context learning and prompt engineering, in numerous application settings [74, 75]. A recent study demonstrates the potential of generative LLMs in filtering vulnerability data by identifying and removing low-quality instances using carefully designed prompts [76]. We posit that our work can also be used to curate high-utility, representative examples for prompts or in-context learning inputs, thereby enhancing the effectiveness of these non-training methods.